OpenStudy (anonymous):

Find the X-intercept of the graph of the equation. 2x+5y=14

OpenStudy (anonymous):

If you want the X-intercept, you want the place on the graph where Y = 0, where the line crosses the Y axis. 2x + 5y = 14 Change Y to zero 2x + 5(0) = 14 2x + 0 = 14 2x = 14 What does X equal? The X-intercept is at ( __, 0 )

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

2X = 14 what does X equal?

OpenStudy (anonymous):

divide both sides of the equation by 2 2X = 14 2X / 2 = 14 / 2 X = ___